OpenStudy (anonymous):

help! Given f(x) = the quantity of 5x plus 1, divided by 2, solve for f-1(8).

OpenStudy (xapproachesinfinity):

f(x)=(5x+1)/2 first find the inverse function, Do you know how to

OpenStudy (anonymous):

no

OpenStudy (xapproachesinfinity):

okay, let me explain it then

OpenStudy (xapproachesinfinity):

First you check if the function passes the horizontal line test. in this case it passes the test since that's the equation of a line

OpenStudy (anonymous):

okay

OpenStudy (xapproachesinfinity):

then Do this y=(5x+1)/2 switch x and y you will get x=(5y+1)/2 then solve for y 2x=5y+1 2x-1=5y (2x-1)/5=y now instead of y replace it with f^-1(x) so f-1(x)=(2x-1)/5

OpenStudy (xapproachesinfinity):

now you can evaluate for f-1(8) by replacing x with 8

OpenStudy (anonymous):

so it would be f-1(8)=2(8)-1/5

OpenStudy (xapproachesinfinity):

correct!

OpenStudy (anonymous):

I got 79/5

OpenStudy (xapproachesinfinity):

No, because you doing this 2(8) -1/5

OpenStudy (xapproachesinfinity):

(2(8)-1)/5 the whole thing is divided by 5 not only 1

OpenStudy (xapproachesinfinity):

If you pay attention i was using parenthesis so you won't confuse that

OpenStudy (anonymous):

oh okay so the answer is 3

OpenStudy (xapproachesinfinity):